# Online Estimation with Rolling Validation: Adaptive Nonparametric Estimation with Streaming Data

### v3

@@ 1 Introduction
We will use $\mathbb{R}$ ($\mathbb{R}^{+}$) to denote (positive) real numbers and $\mathbb{N}^{+}$ to denote the set of positive integers. For $n\in\mathbb{N}^{+}$, we denote $\{1,2,...,n\}$ as $[n]$. For any vector $x$ in $\mathbb{R}^{p}$ or $(\mathbb{N}^{+})^{p}$, we use $x[j]$ to denote its $j$-th element. Notation $a\vee b=\max\{a,b\}$ for $a,b\in\mathbb{R}$. For a random variable $X$ and a probability distribution $P$, the notation $X\sim P$ means “the distribution of $X$ is $P$”. In some conditional expectation notation such as $E[f(X,Y)|F^{i}]$, we use $F^{i}$ to denote the $\sigma$-algebra generated by sample $\{(X_{j},Y_{j}),j\in[i]\}$. The notation $E_{X_{0}}[\hat{f}(X_{0})]$ means taking a conditional expectation with respect to $X_{0}$. We use $\delta_{jk}$ to represent Kronecker delta: $\delta_{jk}=1$ when $j=k$ and $\delta_{jk}=0$ otherwise. The inequality $a\lesssim b,a,b\in\mathbb{R}$ means $a\leq Cb$ for some constant $C$, but the value of $C$ may be different from line to line. When condition $\mathcal{A}$ holds, $\mathbb{1}(\mathcal{A})=1$, otherwise it takes 00.
@@ 2 Stochastic Approximation and the Problem of Online Model Selection
In a typical statistical learning scenario, we have sample points of a pair of random variables $(X,Y)\in\mathbb{R}^{p}\times\mathbb{R}^{1}$ with joint distribution $P$, and want to minimize the predictive risk over a class $\mathcal{F}$ of regression functions: $\min_{f\in\mathcal{F}}E_{(X,Y)\sim P}[l(f;X,Y)]$ under a loss function $l(\cdot)$. We denote the marginal distribution of $X$ as $P_{X}$ and assume that $E(Y^{2})<\infty$. For regression problems, we consider the squared loss $l(f;X,Y)=(f(X)-Y)^{2}$. When $\mathcal{F}$ is the function space of mean-square integrable functions $L_{2}(P_{X})$, the corresponding minimizer
$f_{0}=\operatorname*{argmin}_{f\in L_{2}(P_{X})}E[(f(X)-Y)^{2}]$
@@
In contrast to batch learning where all the data is available at once, streaming samples are usually revealed to the researchers continuously. For computation convenience, it is also increasingly more common to treat large data sets as a data stream and apply online methods by incrementally processing the samples. In both cases, the estimates need to be frequently updated whenever new samples arrive. When a new sample $(X_{i+1},Y_{i+1})\sim P(X,Y)$ arrives, the algorithm uses this new piece of information to update the current estimate $\hat{f}_{i}$ to a new estimate $\hat{f}_{i+1}$. One may convert any batch estimator to an online one by storing all the past data and refitting the whole model repeatedly. But this is often computationally infeasible. For example, fitting a kernel ridge regression estimator using $i$ samples (by matrix inversion) requires $i^{3}$-order computation. This would result in a cumulative $n^{4}$ computation to process $n$ samples in an online fashion. Moreover, batch estimators often need to load the whole data set into the memory, causing significant storage costs, whereas genuine online algorithms only need to load a small portion of data and have a much lower memory footprint.
@@
    The estimate update only depends on the current estimate $\hat{f}_{i}$ and the current data point $(X_{i+1},Y_{i+1})$.
@@
    Each data point is used only once. The past estimates $\{\hat{f}_{j}:j\in[i]\}$ and data points $\{(X_{j},Y_{j}):j\in[i]\}$ are not stored.
@@ 2.1 Examples of Online Estimators and their Hyperparameters
In this section, we review the basic ideas of stochastic approximation, SGD, and the explicit forms of several hyperparameter sequences. Consider a linear model $\mathcal{F}_{\rm linear}$: for each $f\in\mathcal{F}_{\rm linear}$, there exists a vector $\beta\in\mathbb{R}^{p}$ such that $f(X;\beta)=X^{\top}\beta$. Then we arrive at a parametric stochastic optimization problem:
$\min_{\beta\in\mathbb{R}^{p}}E\left[l(f(\cdot;\beta);X,Y)\right].$ (2)
The gradient descent method, a basic numerical optimization technique, uses the following iterative estimation procedure: First, initialize $\tilde{\beta}_{0}=0$, then iteratively update the estimate against the direction of local gradient:
$\tilde{\beta}_{i+1}\leftarrow\tilde{\beta}_{i}-\gamma_{i}\nabla_{\beta}E\left[l(f(\cdot;\beta);X,Y)\right]\big{|}_{\tilde{\beta}_{i}},$
where $\gamma_{i}\in\mathbb{R}^{+}$ is a pre-specified sequence of step sizes (or learning rates). In practice the expectation $E[\cdot]$ cannot be evaluated and must be approximated from the data. SGD procedures replace it with a one-sample (or mini-batch) estimate. In the case of linear regression:
@@
Since $\tilde{f}_{i}(X)=X^{\top}\tilde{\beta}_{i}$ is uniquely determined by the coefficients, we have our first concrete example of the stochastic approximation rule (1). Here the hyperparameter sequence is a sequence of real numbers $\lambda_{i}=\gamma_{i}\in\mathbb{R}^{+}$. Alternatively, we can also re-write the above iteration in a function-updating style:
$\tilde{f}_{i+1}(\cdot)\leftarrow\tilde{f}_{i}(\cdot)+2\gamma_{i}\left(Y_{i+1}-\tilde{f}_{i}\left(X_{i+1}\right)\right)\sum_{k=1}^{p}\phi_{k}\left(X_{i+1}\right)\phi_{k}(\cdot).$ (3)
The function $\phi_{k}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ maps a $p$-dimensional vector to its $k$-th entry $\phi_{k}(X)=X[k]$. We write in this more intricate format to make the above updating rule more comparable with the upcoming nonparametric ones. The Polyak-averaging of $\{\tilde{f}_{i}\}$:
@@
Now we turn to nonparametric online estimation with a diverging model capacity. First, let’s consider the reproducing kernel SGD estimator (kernel-SGD). Let $\mathcal{K}(\cdot,\cdot):\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be a reproducing kernel, with eigen-value/eigen-function pairs $\{(q_{k},\phi_{k}):k\in\mathbb{N}^{+}\}$. The kernel-SGD update rule is
@@
This is our second example of the abstract updating rule (1). The related rate-optimal estimator [10] is its Polyak-averaging $\hat{f}_{i}=i^{-1}\sum_{m=1}^{i}\tilde{f}_{m}$, which can also be updated as described earlier (4).
The performance of kernel-SGD depends on the chosen kernel function $\mathcal{K}$ and the learning rate sequence $\gamma_{i}$. It is shown that a sequence of slowly decreasing step sizes $\gamma_{i}=Ai^{-\zeta}$ would lead to rate-optimal estimators, where $A\in\mathbb{R}^{+}$ and $\zeta\in(0,1/2)$. However, the best choice of $\zeta$ depends on the expansion coefficients $\langle f_{0},\phi_{k}\rangle_{L_{2}(P_{X})}$ and its relative magnitude to the eigenvalues $q_{k}$, which is in general not available to the algorithm. Data-adaptive procedures of choosing $\gamma_{i}$ is needed for better performance of the kernel-SGD in practice. Our framework also allows a varying reproducing kernel such as Gaussian kernels with a varying bandwidth. In the language of (1), the hyperparameters $\lambda_{i}=\gamma_{i}$ is of one dimension when assuming the kernel function is pre-specified and is of two dimension $\lambda_{i}=(\gamma_{i},\sigma_{i})^{\top}$ when there is a varying kernel bandwidth $\sigma_{i}$.
The second line of the kernel-SGD update (5) also suggests a direct way to construct the update from basis expansions without specifying a kernel. Recently, [39, 9] consider “sieve-type” nonparametric online estimators that explicitly use an increasing number of orthonormal basis functions $\{\phi_{k}:k\geq 1\}$, with updating rule
@@
and the averaged version $\hat{f}_{i}$ is similarly updated as in (4). Here $\gamma_{i}$ is a learning rate sequence as before, $J_{i}$ is a positive integer indicating the number of basis functions used in the $(i+1)$-th update, and $\omega$ is a shrinkage parameter which in practice can be taken as $1/2$. Sieve-SGD is computationally more efficient since $J_{i}$ is usually much smaller than $n$ for moderate-dimension problems (where we do not model $p$ increasing with $n$). Under the assumption that $f_{0}$ belong to some Sobolev ellipsoid $W(s)$ spanned by $\{\phi_{k}\}$:
$W(s)=W(s;Q)=\left\{f=\sum_{k=1}^{\infty}\beta_{k}\phi_{k}\text{ such that }\sum_{k=1}^{\infty}\left(\beta_{k}k^{s}\right)^{2}\leq Q^{2}\right\}.$ (7)
In [39], the authors showed that the sieve-SGD has minimal space expense (computer memory cost) among all statistically near-optimal estimators. The theoretical optimal hyperparameters are
$(\gamma_{i},J_{i})=(Ai^{-(2s+1)^{-1}},Bi^{(2s+1)^{-1}}),$ (8)
where $A,B\in\mathbb{R}^{+}$ and $s\geq 1/2$. The hyperparameter sequence $\{\lambda_{i}\}$ of sieve-SGD (6) is typically of two-dimension: $\lambda_{i}=(\gamma_{i},J_{i})^{\top}$.
@@ 2.2 Model Selection in Batch and Online Settings
Suppose we have a data set of size $i$ (or have collected $i$ samples from a data stream) and $K$ estimates $\hat{f}_{i}^{(k)}$, $k\in[K]$. One of the main goals for model selection is identifying $k_{i}^{*}$ such that
$k_{i}^{*}=\operatorname*{argmin}_{k\in[K]}E[r_{i,k}]$ (9)
where
$r(\hat{f})$ $=E_{X_{0}}[\hat{f}(X_{0})-f_{0}(X_{0})]^{2}$ (10)
$r_{i,k}$ $=r(\hat{f}_{i}^{(k)})$
Here $X_{0}$ is a random variable independent from the estimator $\hat{f}$ and $r(\hat{f})$, the risk of $\hat{f}$, still possesses the randomness from constructing $\hat{f}$. The expectation in Equation 9 is marginal, making $E[r_{i,k}]$ a real number only depends on $(i,k)$.
@@
Another way to formulate the model selection problem with an infinite sample size horizon is asymptotically identifying a $k^{*}\in[K]$ —if it exists—such that
$\limsup_{i\rightarrow\infty}\ E[r_{i,k^{*}}]/E[r_{i,k}]<1,\text{ for all }k\in[K]\backslash\{k^{*}\}\,.$ (11)
@@
For instance, in the sieve-SGD example considered in Section 2.1, suppose we only need to choose $\{J_{i}\}$, the number of bases to be used in the estimator. Existing theory [39, 9] suggests that the best choice of $J_{i}$ has the form $J_{i}=Bi^{-s}$ for a pair of constants $(B,s)$. Thus the candidate set of models corresponds to a collection of pairs $\{(B^{(1)},s^{(1)}),...,(B^{(K)},s^{(K)})\}$, which corresponds to $K$ infinite sequences of basis function numbers
$\mathbf{J}^{(k)}=\left\{J_{i}^{(k)}=B^{(k)}i^{\left(2s^{(k)}+1\right)^{-1}},i\in\mathbb{N}\right\}\,,~{}~{}k\in[K].$ (12)
We would, correspondingly, obtain $K$ sequences of estimators $\{\hat{f}^{(k)}_{i},i\in\mathbb{N}\},k\in[K]$, each of them is trained with a different sequence of $\{J_{i}^{(k)}\}$. An online model selection procedure should point us to a choice of $\mathbf{J}_{i}^{(k)}$ that will eventually become the best for each $i$, which will further imply which $B$ and $s$ combination is the best based on available data. In contrast, the standard batch model selection methods only focus on a single sample size and do not suggest a good choice of model for a different sample size. For example, if batch model selection returns a choice of $J=10$ at sample size $100$, it is unclear whether this $J$ comes from a $(B,s)$ combination of $(1,~{}1/2)$ or $(\sqrt{10},~{}1/4)$, which would lead to different choices of $J$ for other values of sample size.

### v4

@@ 1 Introduction
We will use $\mathbb{R}$ ($\mathbb{R}^{+}$) to denote (positive) real numbers and $\mathbb{N}^{+}$ to denote the set of positive integers. For $n\in\mathbb{N}^{+}$, we denote $\{1,2,...,n\}$ as $[n]$. For any vector $x$ in $\mathbb{R}^{p}$ or $(\mathbb{N}^{+})^{p}$, we use $x[j]$ to denote its $j$-th element. Notation $a\vee b=\max\{a,b\}$ for $a,b\in\mathbb{R}$. In some conditional expectation notation such as $\mathbb{E}[f(X,Y)|F^{i}]$, we use $F^{i}$ to denote the $\sigma$-algebra generated by sample $\{(X_{j},Y_{j}),j\in[i]\}$. The notation $\mathbb{E}_{X_{0}}[\hat{f}(X_{0})]$ means taking a conditional expectation with respect to $X_{0}$. We use $\delta_{jk}$ to represent Kronecker delta: $\delta_{jk}=1$ when $j=k$ and $\delta_{jk}=0$ otherwise. The inequality $a\lesssim b,a,b\in\mathbb{R}$ means $a\leq Cb$ for some constant $C$, but the value of $C$ may be different from line to line. When condition $\mathcal{A}$ holds, $\mathbb{1}(\mathcal{A})=1$, otherwise it takes 00.
@@ 2 Stochastic Approximation and the Problem of Online Model Selection
In a typical statistical learning scenario, we have sample points of a pair of random variables $(X,Y)\in\mathbb{R}^{p}\times\mathbb{R}^{1}$ with joint distribution $P$, and want to minimize the predictive risk over a class $\mathcal{F}$ of regression functions: $\min_{f\in\mathcal{F}}\mathbb{E}_{(X,Y)\sim P}[l(f;X,Y)]$ under a loss function $l(\cdot)$. We denote the marginal distribution of $X$ as $P_{X}$ and assume that $\mathbb{E}[Y^{2}]<\infty$. For regression problems, we consider the squared loss $l(f;X,Y)=(f(X)-Y)^{2}$. When $\mathcal{F}$ is the function space of square-integrable functions $L_{2}(P_{X})$, the corresponding minimizer
$f_{0}=\operatorname*{argmin}_{f\in L_{2}(P_{X})}\mathbb{E}[(f(X)-Y)^{2}]$
@@
In contrast to batch learning, where all the data is available at once, streaming samples are usually revealed to the researchers continuously. For computational convenience, it is also increasingly more common to treat large data sets as a data stream and apply online methods by incrementally processing the samples. In both cases, the estimates need to be frequently updated whenever new samples arrive. When a new sample $(X_{i+1},Y_{i+1})\sim P(X,Y)$ arrives, the algorithm uses this new piece of information to update the current estimate $\hat{f}_{i}$ to a new estimate $\hat{f}_{i+1}$. One may convert any batch estimator to an online one by storing all the past data and refitting the whole model repeatedly. But this is often computationally infeasible. For example, fitting a kernel ridge regression estimator using $i$ samples (by matrix inversion) requires $i^{3}$-order computation. This would result in a cumulative $n^{4}$ computation to process $n$ samples in an online fashion. Moreover, batch estimators often need to load the whole data set into memory, causing significant storage costs, whereas genuine online algorithms operate incrementally and maintain a much smaller memory footprint.
@@
    The estimate update only depends on the current estimate $\hat{f}_{i}$ and the current data point.
@@
    Each data point is used only once. The past estimates $\{\hat{f}_{j}:j\in[i-1]\}$ and data points $\{(X_{j},Y_{j}):j\in[i-1]\}$ are not stored.
@@ 2.1 Examples of Online Estimators and their Hyperparameters
In this section, we review the basic ideas of stochastic approximation, SGD, and the explicit forms of several hyperparameter sequences. Consider a linear model $\mathcal{F}_{\rm linear}$: for each $f\in\mathcal{F}_{\rm linear}$, there exists a vector $\beta\in\mathbb{R}^{p}$ such that $f(X;\beta)=X^{\top}\beta$. This brings us to a parametric stochastic approximation problem:
$\min_{\beta\in\mathbb{R}^{p}}\mathbb{E}[l(f(\ \cdot\ ;\beta);X,Y)].$
The gradient descent method, a basic numerical optimization technique, uses the following iterative estimation procedure: First, initialize $\tilde{\beta}_{0}=0$, then iteratively update the estimate against the direction of the local gradient:
$\tilde{\beta}_{i+1}\leftarrow\tilde{\beta}_{i}-\gamma_{i}\nabla_{\beta}\mathbb{E}\left[l(f(\ \cdot\ ;\beta);X,Y)\right]\big{|}_{\tilde{\beta}_{i}},$
where $\gamma_{i}\in\mathbb{R}^{+}$ is a pre-specified sequence of step sizes (or learning rates). In practice, the expectation $\mathbb{E}[\ \cdot\ ]$ cannot be evaluated and must be approximated from the data. SGD procedures replace it with a one-sample (or mini-batch) estimate. In the case of linear regression:
@@
We have our first concrete example of the stochastic approximation rule (1). Here, the hyperparameter sequence is a real number series $\lambda_{i}=\gamma_{i}\in\mathbb{R}^{+}$. Alternatively, we can rewrite the above iteration in a function-updating style:
$\tilde{f}_{i+1}(\cdot)\leftarrow\tilde{f}_{i}(\cdot)+2\gamma_{i}\left(Y_{i+1}-\tilde{f}_{i}\left(X_{i+1}\right)\right)\sum_{k=1}^{p}\phi_{k}\left(X_{i+1}\right)\phi_{k}(\cdot).$
The function $\phi_{k}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ maps a $p$-dimensional vector to its $k$-th entry $\phi_{k}(X)=X[k]$. We present it in this more intricate format to better align with the nonparametric versions that follow. The Polyak-averaging of $\{\tilde{f}_{i}\}$:
@@
Now we turn to nonparametric online estimation with a diverging model capacity. First, let’s consider the reproducing kernel SGD estimator (kernel-SGD). Let $\mathcal{K}(\cdot,\cdot):\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be a reproducing kernel with associated eigenvalue–eigenfunction pairs $\{(q_{k},\phi_{k}):k\in\mathbb{N}^{+}\}$. The kernel-SGD update rule is
@@
This is our second example of the abstract update rule (1). The related rate-optimal estimator [10] is its Polyak-averaging $\hat{f}_{i}=i^{-1}\sum_{m=1}^{i}\tilde{f}_{m}$, which can also be updated as described earlier (2).
The performance of kernel-SGD depends on the selected kernel function $\mathcal{K}$ and the learning rate $\gamma_{i}$. It has been shown that using a sequence of slowly decaying step sizes $\gamma_{i}=Ai^{-\zeta}$ yields rate-optimal estimators, where $A\in\mathbb{R}^{+}$ and $\zeta\in(0,1/2)$. However, the best choice of $\zeta$ depends on the expansion coefficients $\langle f_{0},\phi_{k}\rangle_{L_{2}(P_{X})}$ and its relative magnitude to the eigenvalues $q_{k}$, which is in general not available to the algorithm. Data-adaptive procedures for choosing $\gamma_{i}$ are needed for better performance of the kernel-SGD in practice. Our framework also allows a varying reproducing kernel, such as Gaussian kernels with a dynamic bandwidth. In the notation of (1), the hyperparameter $\lambda_{i}=\gamma_{i}$ is one-dimensional when the kernel function is fixed, and becomes two-dimensional, $\lambda_{i}=(\gamma_{i},\sigma_{i})^{\top}$, when the kernel bandwidth $\sigma_{i}$ is allowed to vary.
The second line of the kernel-SGD update (3) also suggests a direct way to construct the update from basis expansions without specifying a kernel. Recently, [40, 9] consider “sieve-type” nonparametric online estimators that explicitly use an increasing number of orthonormal basis functions $\{\phi_{k}:k\geq 1\}$, with an update rule
@@
and the averaged version $\hat{f}_{i}$ is similarly updated as in (2). Here, $\gamma_{i}$ denotes the learning rate sequence as before, $J_{i}$ is a positive integer representing the number of basis functions used in the $(i+1)$-th update, and $\omega$ is a shrinkage parameter, which can typically be set to $1/2$ in practice. Sieve-SGD (4) is computationally more efficient since $J_{i}$ is usually much smaller than $n$ for moderate-dimension problems (where we do not model $p$ increasing with $n$). Under the assumption that $f_{0}$ belong to some Sobolev ellipsoid $W(s)$ spanned by $\{\phi_{k}\}$:
$W(s)=W(s;Q)=\left\{f=\sum_{k=1}^{\infty}\beta_{k}\phi_{k}\text{ such that }\sum_{k=1}^{\infty}\left(\beta_{k}k^{s}\right)^{2}\leq Q^{2}\right\},$ (5)
it has been shown [40] that the sieve-SGD has minimal space expense (computer memory cost) among all statistically optimal estimators. The theoretical optimal hyperparameters are
$\left(\gamma_{i},J_{i}\right)=\left(Ai^{-1/(2s+1)},Bi^{1/(2s+1)}\right),$ (6)
where $A,B\in\mathbb{R}^{+}$ and $s\geq 1/2$. The hyperparameter sequence $\{\lambda_{i}\}$ of sieve-SGD (4) is typically two-dimensional, with $\lambda_{i}=(\gamma_{i},J_{i})^{\top}$.
@@ 2.2 Model Selection in Batch and Online Settings
Suppose we have a dataset of size $i$ (or have collected $i$ samples from a data stream), along with $K$ estimates $\hat{f}_{i}^{(k)}$ for $k\in[K]$. One of the main goals of model selection is identifying the optimal index $k_{i}^{*}$ that minimizes the expected risk:
$k_{i}^{*}=\operatorname*{argmin}_{k\in[K]}\mathbb{E}[r_{i,k}]$ (7)
where the risk is defined as
$r(\hat{f})$ $=\mathbb{E}_{X_{0}}\left[\left(\hat{f}\left(X_{0}\right)-f_{0}\left(X_{0}\right)\right)^{2}\right]$
$r_{i,k}$ $=r\big{(}\hat{f}_{i}^{(k)}\big{)}.$
Here $X_{0}$ is a random variable independent from the estimator $\hat{f}$. And $r(\hat{f})$, the risk of $\hat{f}$, still possesses the randomness from constructing $\hat{f}$. The expectation in Equation 7 is marginal, making $\mathbb{E}[r_{i,k}]$ a real number that only depends on $(i,k)$.
@@
Another way to formulate the model selection problem in the infinite-horizon setting is to asymptotically identify an index $k^{*}\in[K]$—if it exists—such that
$\limsup_{i\rightarrow\infty}\ \mathbb{E}[r_{i,k^{*}}]/\mathbb{E}[r_{i,k}]<1,\text{ for all }k\in[K]\backslash\{k^{*}\}\,.$ (8)
@@
For instance, in the sieve-SGD example discussed in Section 2.1, suppose the only hyperparameter to be selected is ${J_{i}}$, the number of basis functions used in the estimator. Existing theory [40, 9] suggests that the best choice of $J_{i}$ has the form $J_{i}=Bi^{1/(2s+1)}$ for a pair of constants $(B,s)$. Thus the candidate set of models corresponds to a collection of pairs $\{(B^{(1)},s^{(1)}),...,(B^{(K)},s^{(K)})\}$, each defining one of $K$ infinite sequences specifying the number of basis functions:
$\mathbf{J}^{(k)}=\left\{J_{i}^{(k)}=B^{(k)}i^{1/\left(2s^{(k)}+1\right)},i\in\mathbb{N}\right\}\,,~{}~{}k\in[K].$
Correspondingly, we obtain $K$ sequences of estimators $\{\hat{f}^{(k)}_{i},i\in\mathbb{N}\},k\in[K]$, where each of them is trained with a different $\mathbf{J}^{(k)}$. An online model selection procedure should guide us toward the sequence $\mathbf{J}^{(k)}$ that eventually performs best, thereby identifying the optimal combination of $(B^{(k)},s^{(k)})$ based on the observed data. In contrast, batch model selection methods focus more on a single sample size and do not suggest a good choice of model for a different sample size. For example, if batch model selection returns a choice of $J=10$ at sample size $100$, it is unclear whether this $J$ comes from a $(B,s)$ combination of $(1,~{}1/2)$ or $(\sqrt{10},~{}3/2)$, which would lead to different choices of $J$ for other values of sample size.